# SAVA: Scalable Learning-Agnostic Data Valuation

### v1

@@ 3 SAVA: Scalable Data Valuation
We repeat this process and solve the OT problem for each pair of $(B_{i},B^{\prime}_{j})$, i.e., OT problem for distributions on batches of data points, across the training and validation datasets, i.e., for all $i\in[K_{t}],j\in[K_{v}]$. This enables us to define the cost matrix for pairwise batches in $\bar{\mu}_{t},\bar{\mu}_{v}$, denoted as ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}(\bar{\mu}_{t},\bar{\mu}_{v})\in{\mathbb{R}}^{K_{t}\times K_{v}}_{+}$ where we recall that $K_{t}$ and $K_{v}$ are the number of batches in training and validation sets. Hence, for this cost matrix ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$, the element ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}_{ij}$ is computed as $\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, for all $i\in[K_{t}],\forall j\in[K_{v}]$.
Batch valuation. Given the pairwise cost matrix across batches ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$, we compute the data valuation for each batch via the gradient of the distance $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ w.r.t. the probability mass of batches in the two datasets, i.e., $\frac{\partial\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\partial\bar{\mu}_{t}({B}_{i})}$. These partial derivatives measure the contribution of the batches to the OT distance, i.e., shifting more probability mass to the batch would result in an increase or decrease of the dataset distance, respectively.
@@
Since the optimal dual variables are only unique up to a constant, we follow Just et al., (2023) to normalize these optimal dual variables such that the sum of all elements is equal to a zero vector:
@@
Using batch valuation for data point valuation. After solving the OT problem over batches, we obtain the optimal transport plan ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ which is used to compute the data evaluation over individual data points:
${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}=\text{diag}{(\bar{f}_{k}^{*})}\exp\Bigl{(}-\frac{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\epsilon}\Bigr{)}\text{diag}{(\bar{g}_{k}^{*})},$ (11)
where $\text{diag}(\cdot)$ is matrix diagonal operator.
@@
Using HOT, we can measure the gradients of the OT  distance w.r.t. the probability mass of a given data point in each dataset via the calibrated gradient summarized in the following Lemma 1. We refer to Appendix Table 1 for the notations.
@@
The data valuation, i.e., the calibrated gradient for a data point $z_{k}$ in the batch ${B}_{i}$ in ${\mathbb{D}}_{t}$ can be computed as follows:
@@
As the common practice in computing the OT via its entropic regularization, using the Sinkhorn algorithm (Cuturi,, 2013), we quantify the deviation in the calibrated gradients caused by the entropy regularizer. This analysis provides foundations on the potential impact of the deviation on the applications built on these gradients.
@@
Let $\textnormal{HOT}(\mu_{t},\mu_{v})$ and $\textnormal{OT}_{\epsilon}(\mu_{t},\mu_{v})$ be the HOT and entropic regularized OT between the measures $\mu_{t}$ and $\mu_{v}$ associated with the two datasets ${\mathbb{D}}_{t}$ and ${\mathbb{D}}_{v}$ respectively. The difference between the calibrated gradients for two data points $\{z_{l},z_{h}\}\in B_{i}\subset{\mathbb{D}}_{t}$ can be calculated as
$\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}-\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{l})}=\sum_{j=1}^{K_{v}}{\color[rgb]{1,.5,0}\bar{\pi}^{*}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}\Bigl{[}\frac{\partial\textnormal{OT}_{\epsilon}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})}-\frac{\partial\textnormal{OT}_{\epsilon}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{l})}$
$-\epsilon\frac{N_{k}}{N_{k}-1}\left(\frac{1}{(\bar{\pi}^{*}_{\epsilon})_{l,j}}-\frac{1}{(\bar{\pi}^{*}_{\epsilon})_{k,j}}\right)\Bigr{]}.$ (15)
@@
Refer to Appendix C for the Proof. ∎
We make a similar observation in LAVA that the ground-truth gradient difference between two training points $z_{k}$ and $z_{l}$ is calculated based on the HOT formulation and can be approximated by the entropic regularized formulation $\textnormal{OT}_{\epsilon}$, such as via Sinkhorn algorithm (Cuturi,, 2013). In other words, we can calculate the ground-truth difference based on the solutions to the regularized problem plus some calibration terms that scale with $\epsilon$. In addition, in our case with HOT, the gradient difference also depends on the additional optimal assignment across batches ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ which is again estimated by Sinkhorn algorithm.
@@ 4 Properties and Discussions
We then solve the OT problem across batches $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ to get the optimal plan ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ in line 6 Algorithm 1. The cost matrix is of size $K_{t}\times K_{v}$ which is typically small and so less expensive compared to the previous OT tasks. Finally, from Lines 7 to 10, we estimate valuation scores for training data using the auxiliary matrices computed in the previous steps, including ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}$ and ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$.
@@
Practical Implementation with Caching. While yielding a significant memory improvement, SAVA’s runtime is not necessarily faster than LAVA. To speed up SAVA, we propose to implement Algorithm 1 by caching the label-to-label costs between points in the validation and training batches: $\textnormal{OT}_{{\textnormal{d}}}(\cdot,\cdot)$ in Eq. (1) so that it is only calculated once in the first iteration of Algorithm 1 and reused for subsequent batches. This significantly reduces SAVA runtimes with no detriment to performance Figure 9. All experimental results in Section 5, unless otherwise stated, implement this caching strategy.
Batch Sizes. If we consider Eq. (7), HOT provides the upper bound for the OT since its optimal coupling is on batches of data points. HOT recovers the OT when either batch only has one support or each dataset only has one batch. Consequently, up to a certain batch size, when increasing the batch size $N_{i}$, HOT converges to the OT, but its memory complexity also increases, i.e., $\mathcal{O}(N^{2}_{i})\rightarrow\mathcal{O}(N^{2})$. On the other hand, if the batch size $N_{i}$ is too small, the number of batches $K_{t}$ will be large. As a result, the memory complexity will also be high (i.e., for solving $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ in Algorithm 1). Thus, the batch size will trade off the memory complexity and the approximation of HOT to OT.
5 Experiments
We aim to test two following hypotheses [^1] The code to reproduce all experiments can be found at https://github.com/skezle/sava.:
1. Can SAVA scale and overcome the memory complexity issues which hinder LAVA?
2. Can SAVA scale to a large real-world noisy dataset with over a million data points?
@@ 5.1 Dataset Corruption Detection
We test the scalability of SAVA versus LAVA (Just et al.,, 2023) by leveraging the CIFAR10 dataset, introducing a corruption to a percentage of the training data, but keeping the validation set clean. We then assign a value to each data point in the training set. Following Pruthi et al., (2020); Just et al., (2023), we sort the training examples by their values. An effective data valuation method would rank corrupted examples in a prefix of ordered points. We use the fraction of corrupted data recovered by the prefix of size $N/4$ as our detection rate (see an example in Appendix EFigure 6).
[…]
@@
Setup. We consider $4$ different corruptions (see Appendix D for details) that are applied to the training data following the settings in Just et al., (2023): (i) noisy labels, (ii) noisy features, (iii) backdoor attacks and (iv) poison detections. All experiments are run on a Tesla K80808080 Nvidia GPUs with $12$GB GPU RAM. Unless otherwise stated, all results reported are a mean ±$\pm 1$ standard deviation over $5$ independent runs.
@@
Baselines. Our main baseline is LAVA which shares the same design scheme based on hierarchically defined OT for data valuation Eq. (1). For SAVA and LAVA we use features from a pre-trained ResNet18 (He et al.,, 2016). For SAVA, we use a default batch size of $N_{i}=1024$ which is its main hyperparameter. [^2] See Appendix G for the study of the sensitivity of the batch size $N_{i}$ in SAVA. We also consider KNN Shapley (Jia et al., 2019a, ), the Shapley value measures the marginal improvement in the utility of a data point and uses KNN to approximate the Shapley value. KNN Shapley also uses ResNet18 features to calculate Shapley values, we tune $k$ as its performance is sensitive to this hyperparameter. We also consider TracInCP (Pruthi et al.,, 2020) which measures the influence of each training point by measuring the difference in the loss at the beginning versus the end of training (see Section F.1 for details). We also compare with Influence Function (Koh and Liang,, 2017) which approximates the effect of holding out a training point on the test error, it uses expensive approximations of the Hessian to calculate the influence of a training point. We finally consider a naive OT baseline which obtains values for data points at a batch level, and aggregates values across validation batches; the baseline essentially performs LAVA only at a data point level within a batch without a batch level OT problem, line 6 Algorithm 1. We call this baseline Batch-wise LAVA. Although Batch-wise LAVA obtains good corruption detection results, it is very sensitive to the batch size (see Appendix G).
@@
Noisy labels. We corrupt $30\%$ of the labels in the training set by randomly assigning the target a different label. From Figure 2, we can see that LAVA has an out-of-memory (OOM) issue when valuing the full training set of $50$K points. In contrast, SAVA, Batch-wise LAVA KNN Shapley can consistently value and detect all corruptions when inspecting $12.5$K ordered samples by values. Influence Functions matches the performance of SAVA, but is very expensive to run on large datasets. TracInCP is unable to detect noisy labels better than random selection, similar to the observations in Just et al., (2023). Consequently, when pruning $30\%$ of the data, SAVA can consistently improve its accuracy on the test set as the training set size increases (Figure 3).
@@
Noisy features. We add Gaussian noise to $30\%$ of the training images to simulate feature corruptions that might occur in real datasets. From Figure 2, LAVA obtains good performance for moderate dataset sizes, but for training set sizes above $10$K, some runs have OOM errors when embedding the entire training set into memory and when calculating the full cost matrix for OT problem in Eq. (1). In contrast, SAVA can consistently value and detect corruptions. Batch-wise LAVA is also able to scale and gets similar performance to SAVA. As can KNN Shapley, albeit its detection rate is lower than SAVA. TracInCP and Influence Functions both struggle to detect noisy data points, both methods were originally shown to detect noisy labels, so we do not expect them to work well beyond noisy label detection.
@@
Backdoor attacks. We corrupt $10\%$ of the data with a Trojan square attack (Liu et al.,, 2018). SAVA, Batch-wise LAVA and KNN Shapley can scale as the dataset size increases while LAVA has an OOM error when valuing the largest dataset with $50$k points. TracInCP and Influence Functions both struggle to detect backdoored training points and have long runtimes.
@@
Poisoning attacks. We corrupt $10\%$ of the data with a poison frogs attack (Shafahi et al.,, 2018). We find in Figure 2 that SAVA and Batch-wise LAVA can scale and maintain a high detection rate. In contrast, KNN Shapley struggles to detect many corrupted data points after inspecting $25\%$ of the ordered training data. Both TracInCP and Influence Functions struggle to detect the corrupted data.
Pruning. We can prune the $N/4$ data points and train a classifier on the pruned dataset in Figure 3. SAVA can value a larger and larger pool of training data. The subsequently pruned dataset provides better and better performance as more clean data - which resembles the validation set - is used for training. In contrast, LAVA has an OOM issue for the largest dataset when running the Sinkhorn algorithm on a training set of size $50$K. As a result, the performance of the ResNet18 model does not improve when valuing larger training sets with LAVA.

### v2

@@ 3 SAVA: Scalable Data Valuation
We repeat this process and solve the OT problem for each pair $(B_{i},B^{\prime}_{j})$, i.e., the OT problem for distributions on batches of data points, across the training and validation datasets, for all $i\in[K_{t}],j\in[K_{v}]$. This enables us to define the cost matrix for pairwise batches in $\bar{\mu}_{t},\bar{\mu}_{v}$, denoted as ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}(\bar{\mu}_{t},\bar{\mu}_{v})\in{\mathbb{R}}^{K_{t}\times K_{v}}_{+}$ where we recall that $K_{t}$ and $K_{v}$ are the number of batches in training and validation sets. Hence, for this cost matrix ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$, the element ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}_{ij}$ is computed as $\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, for all $i\in[K_{t}],\forall j\in[K_{v}]$.
Batch valuation. Given the pairwise cost matrix across batches ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$, we compute the data valuation for each batch via the (sub)gradient of the distance $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ w.r.t. the probability mass of batches in the two datasets, i.e., $\frac{\partial\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\partial\bar{\mu}_{t}({B}_{i})}$. These partial derivatives measure the contribution of the batches to the OT distance, i.e., shifting more probability mass to the batch would result in an increase or decrease of the dataset distance, respectively.
@@
Since the optimal dual variables are only unique up to a constant, we follow just2023lava to normalize these optimal dual variables such that the sum of all elements is equal to a zero vector:
@@
Using batch valuation for data point valuation. After solving the OT problem over batches, we obtain the optimal transport plan ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ which is used to compute the data valuation over individual data points:
${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}=\text{diag}{(\bar{u}^{*})}\exp\Bigl{(}-\frac{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\varepsilon}\Bigr{)}\text{diag}{(\bar{v}^{*})},$ (10)
where $\text{diag}(\cdot)$ is matrix diagonal operator, $(\bar{u}^{*},\bar{v}^{*})$ is a dual solution from Sinkhorn algorithm. Additionally, we have $\bar{u}^{*}=\exp(-\frac{1}{2}-\frac{\bar{f}^{*}}{\varepsilon}),\bar{v}^{*}=\exp(-\frac{1}{2}-\frac{\bar{g}^{*}}{\varepsilon})$ following Cuturi-2013-Sinkhorn.
@@
Using HOT, we can measure the (sub)gradients of the OT  distance w.r.t. the probability mass of a given data point in each dataset via the calibrated gradient summarized in the following Lemma 1. We refer to Appendix Table LABEL:table:notation for the notations.
@@
The calibrated gradient for a data point $z_{k}$ in the batch ${B}_{i}$ in ${\mathbb{D}}_{t}$ can be computed as:
@@
It is common practice to compute the OT via its entropic regularization, using the Sinkhorn algorithm (Cuturi-2013-Sinkhorn). We quantify the deviation in the calibrated gradients caused by the entropy regularizer. This analysis shows the potential impact of the deviation on applications built on these gradients.
We refine the theoretical result (just2023lava, Theorem 2), which uses entropic regularization to approximate the OT (sub)gradient in two aspects. Firstly, we take into account the non-negative constraint of the OT plan $\pi\geq 0$ in the dual formulation. Secondly, we correct the discrete formulation for entropic regularization (see LABEL:appendix_issue_theorem2_lava for the discussion).
Theorem 2 (Refined Theorem 2 in just2023lava).
The difference between calibrated gradients for two data points in ${\mathbb{D}}_{t}$ and ${\mathbb{D}}_{v}$ can be calculated as
$\frac{\partial\operatorname{OT}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{i}\right)}-\frac{\partial\operatorname{OT}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{k}\right)}=\frac{\partial\textnormal{OT}_{\varepsilon}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{i}\right)}-\frac{\partial\textnormal{OT}_{\varepsilon}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{k}\right)}$
$\hskip 170.00026pt+\varepsilon\frac{N}{N-1}\left({\color[rgb]{0,0,1}\log\frac{\left(\pi_{\varepsilon}^{*}\right)_{ij}/\mu_{t}(z_{i})}{\left(\pi_{\varepsilon}^{*}\right)_{kj}/\mu_{t}(z_{k})}+\frac{h_{kj}^{*}-h_{ij}^{*}}{\varepsilon}}\right),$ (14)
where $h^{*}$ is the corresponding optimal dual variable, accounting for non-negative constraint on the transportation plan $\pi\geq 0$ in the primal OT formulation.
Proof.
Refer to Appendix LABEL:appendix_proof_theorem2 for the proof. ∎
The term in the brackets (in blue) is the key difference between our new result and the previous one derived in LAVA (just2023lava, Theorem 2).
@@
Let $\textnormal{HOT}(\mu_{t},\mu_{v})$ and $\textnormal{OT}_{\varepsilon}(\mu_{t},\mu_{v})$ be the HOT and entropic regularized OT between the measures $\mu_{t}$ and $\mu_{v}$ associated with the two datasets ${\mathbb{D}}_{t}$ and ${\mathbb{D}}_{v}$ respectively. The difference between the calibrated gradients for two data points $\{z_{l},z_{h}\}\in B_{i}\subset{\mathbb{D}}_{t}$ can be calculated as
$\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}-\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{l})}=\sum_{j=1}^{K_{v}}{\color[rgb]{1,.5,0}\bar{\pi}^{*}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}\Bigl{[}\frac{\partial\textnormal{OT}_{\varepsilon}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})}-\frac{\partial\textnormal{OT}_{\varepsilon}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{l})}$
$\hskip 150.00023pt+\varepsilon\frac{N_{i}}{N_{i}-1}\left(\log\frac{(\bar{\pi}^{*}_{\varepsilon})_{k,j}/\mu_{t}(z_{k})}{(\bar{\pi}^{*}_{\varepsilon})_{l,j}/\mu_{t}(z_{l})}+{\frac{h_{lj}^{*}-h_{kj}^{*}}{\varepsilon}}\right)\Bigr{]}.$ (15)
@@
Refer to Appendix LABEL:appendix_proof_lemma3 for the proof. ∎
We make a similar observation in LAVA that the ground-truth (sub)gradient difference between two training points $z_{k}$ and $z_{l}$ is calculated based on the HOT formulation and can be approximated by the entropic regularized formulation $\textnormal{OT}_{\varepsilon}$, such as via the Sinkhorn algorithm (Cuturi-2013-Sinkhorn). In other words, we can calculate the ground-truth difference based on the solutions to the regularized problem plus some calibration terms that scale with $\varepsilon$. In addition, in our case with HOT, the (sub)gradient difference also depends on the additional optimal assignment across batches ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ which is again estimated by the Sinkhorn algorithm.
@@ 4 Properties and Discussions
In line 6 Algorithm 1, we solve the OT problem across batches $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ to obtain ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$. The cost matrix is small, of size $K_{t}\times K_{v}$, so less expensive compared to working with the full dataset just2023lava. Finally, from Lines 7 to 10, we estimate valuation scores for training data using the auxiliary matrices computed in the previous steps, including ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}$ and ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$.
@@
Practical Implementation with Caching. While yielding a significant memory improvement, SAVA’s runtime is not necessarily faster than LAVA. To speed up SAVA, we propose to implement Algorithm 1 by caching the label-to-label costs between points in the validation and training batches: $\textnormal{OT}_{{\textnormal{d}}}(\cdot,\cdot)$ in Sec. 2.1 so that it is only calculated once in the first iteration of Algorithm 1 and reused for subsequent batches. This significantly reduces SAVA runtimes with no detriment to performance LABEL:fig:l2l_caching_ablation. All experimental results in Section 5, unless otherwise stated, implement this caching strategy.
Batch Sizes. If we consider Eq. (6), HOT provides the upper bound for the OT since its optimal coupling is on batches of data points. HOT recovers the OT when either batch only has one support or each dataset only has one batch. Consequently, up to a certain batch size, when increasing the batch size $N_{i}$, HOT converges to the OT, but its memory complexity also increases, i.e., $\tilde{\mathcal{O}}(N^{2}_{i})\rightarrow\tilde{\mathcal{O}}(N^{2})$. On the other hand, if the batch size $N_{i}$ is too small, the number of batches $K_{t}$ will be large. As a result, the memory complexity will also be high (i.e., for solving $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ in Algorithm 1). Thus, the batch size will trade off the memory complexity and the approximation of HOT to OT.
5 Experiments
We aim to the test two following hypotheses: 1. Can SAVA scale and overcome the memory complexity issues which hinder LAVA while maintaining similar performance? 2. Can SAVA scale to a large real-world noisy dataset with over a million data points?
@@ 5.1 Dataset Corruption Detection
We test the scalability of SAVA versus LAVA (just2023lava) by leveraging the CIFAR10 dataset, introducing a corruption to a percentage of the training data, but keeping the validation set clean. We then assign a value to each data point in the training set. Following pruthi2020estimating; just2023lava, we sort the training examples by their values. An effective data valuation method would rank corrupted examples in a prefix of ordered points. We use the fraction of corrupted data recovered by the prefix of size $N/4$ as our detection rate (see an example in LABEL:app:data_val_rankingsLABEL:fig:data_val_example).
@@
Setup. We consider $4$ different corruptions (see LABEL:appendix_data_corruptions for details) applied to training data following just2023lava: (i) noisy labels, (ii) noisy features, (iii) backdoor attacks and (iv) poison detections. All experiments are run on a Tesla K80808080 Nvidia GPUs with $12$GB GPU RAM. Unless otherwise stated, all results reported are a mean ±$\pm 1$ standard deviation over $5$ independent runs.
@@
Baselines. Our main baseline is LAVA. For SAVA and LAVA we use features from a pre-trained ResNet18 (he2016deep). For SAVA, we use a default batch size of $N_{i}=1024$ which is its main hyperparameter. [^3] See LABEL:sec:ablations for the study of the sensitivity of the batch size $N_{i}$ in SAVA. We consider KNN Shapley (jia2019efficient), the Shapley value measures the marginal improvement in the utility of a data point and uses KNN to approximate the Shapley value. KNN Shapley also uses ResNet18 features to calculate Shapley values, we tune $k$ as its performance is sensitive to this hyperparameter. We consider TracInCP (pruthi2020estimating) which measures the influence of each training point by measuring the difference in the loss at the beginning versus the end of training. We also compare with Influence Function (koh2017understanding) which approximates the effect of holding out a training point on the test error, it uses expensive approximations of the Hessian to calculate the influence of a training point. We also consider Data-OOB (kwon2023data) which uses bagging estimators to value data points, data points are embedded using a pre-trained feature extractor, and the bagging estimator is a decision tree (see LABEL:sec:imp_details_cifar10 for implementation details). We finally consider a naive OT baseline which obtains values for data points at a batch level, and aggregates values across validation batches; the baseline essentially performs LAVA only at a data point level within a batch and averaging values across validation batches. We call this baseline Batch-wise LAVA. Although Batch-wise LAVA obtains good corruption detection results, it is very sensitive to the batch size (see LABEL:sec:batchwise_lava_robustness).
@@
Noisy labels. We corrupt $30\%$ of the labels in the training set by randomly assigning the target a different label. From Figure 2, we can see that LAVA has an out-of-memory (OOM) issue when valuing the full training set of $50$K points. In contrast, SAVA, Batch-wise LAVA, KNN Shapley and Data-OOB can consistently value and detect all corruptions when inspecting $12.5$K ordered samples by values. Influence Functions matches the performance of SAVA, but is very expensive to run on large datasets. TracInCP is unable to detect noisy labels better than random selection, similar to the observations in just2023lava. Consequently, when pruning $30\%$ of the data, SAVA can consistently improve its accuracy on the test set as the training set size increases LABEL:sec:cifar10_pruning.
@@
Noisy features. We add Gaussian noise to $30\%$ of the training images to simulate feature corruptions that might occur in real datasets. From Figure 2, LAVA obtains good performance for moderate dataset sizes, but for training set sizes above $10$K, some runs have OOM errors when embedding the entire training set into memory and when calculating the full cost matrix for OT problem. In contrast, SAVA can consistently value and detect corruptions. Batch-wise LAVA is also able to scale and gets similar performance to SAVA. As can KNN Shapley and Data-OOB, albeit its detection rate is lower than SAVA; as we prune $30\%$ of the data for larger and larger dataset sizes performance of SAVA outperforms KNN Shapley LABEL:fig:cifar10_corruptions_prune. TracInCP and Influence Functions both struggle to detect noisy data points, both methods were originally shown to detect noisy labels, so we do not expect them to work well beyond noisy label detection.
@@
Backdoor attacks. We corrupt $10\%$ of the data with a Trojan square attack (liu2018trojaning). SAVA, Batch-wise LAVA and KNN Shapley can scale as the dataset size increases while LAVA has an OOM error when valuing the largest dataset with $50$k points. TracInCP and Influence Functions both struggle to detect backdoored training points and have long runtimes. Data-OOB also struggles to value data points in this scenario, this is in line with recent surveys which show less strong performance when valuing data with noisy features jiang2023opendataval.
@@
Poisoning attacks. We corrupt $10\%$ of the data with a poison frogs attack (shafahi2018poison). We find in Figure 2 that SAVA and Batch-wise LAVA can scale and maintain a high detection rate. In contrast, KNN Shapley struggles to detect corrupted data points after inspecting $25\%$ of the ordered training data. TracInCP, Data-OOB and Influence Functions struggle to detect the corrupted data.